OpenStudy (anonymous):

The measure of one angle of a triangle is 108 . The measures of the other two angles are 1:5

OpenStudy (anonymous):

180-108= 7272/6+?

OpenStudy (anonymous):

72/6=?

OpenStudy (anonymous):

Huh? 7272?

OpenStudy (anonymous):

I understand the other two angles add up to 72

OpenStudy (anonymous):

What does the ratio mean.

OpenStudy (anonymous):

it means that one angel is 5 times bigger than the other angle

OpenStudy (anonymous):

Ohhh. Thank you,

OpenStudy (anonymous):

i meant to hit enter sorry does that make sense to you?

OpenStudy (anonymous):

Yes. Got it!